Solve: 30x^2+11x-30=0
I'm lost!! Can you help me? Thank you!

Answer by stanbon(75887)

Solve: 30x^2+11x-30=0
Factor to get:
(5x+6)(6x-5)=0
x = -6/5 or x = 5/6
=========================
You could also solve the quadratic with the quadratic formula:
x = [-11+-sqrt(11^2-4*30*-30)]/60
x = [-11+-sqrt3721]/60
x = [-11+-61]/60
x = 50/60 or x = -72/60
x = 5/6 or x= -12/10
x=5/6 or x=-6/5
=====================
Cheers,
Stan H.